your client owns a 14 story building that generates a gross rental revenue of $1,332,900; operating expenses of $412,300; interest expense of$575,000; and incomr taces of $72,000. you are going to lend $9,200,000 at market rate of 6.25% with 25 year amortization. what is the acual debt service cover ratio?

A. 1.17
B. 1.26
C. 1.48
D. 1.6
Answer #1

Net operating income = NOI =  a gross rental revenue of $1,332,900 - operating expenses of $412,300 = 1,332,900 - 412,300 = $ 920,600

Annual mortgage payment = PMT (Rate, Nper, PV, FV) = PMT (6.25%, 25, -9200000, 0) = $ 736,870

Hence, debt service coverage ratio = NOI / Annual mortgage payment = 920,600 / 736,870 = 1.26

Hence, the correct answer is the second option i.e. option B. 1.26